A Chemist has two solutions of alcohol. One has a 24% concentration and the other has a 40% solution. How many liters of EACH solution must be mixed to obtain 120 liters of a 30% solution?
Answer by mananth(16946)

A Chemist has two solutions of alcohol. One has a 24% concentration and the other has a 40% solution. How many liters of EACH solution must be mixed to obtain 120 liters of a 30% solution?
Soln1 soln2 mix
24% 40% 30%
X 120-x 120
0.24x+0.4(120-x)=0.3*120
0.24x+48-0.4x=36
-0.16x=-12
X=12/0.16
X= 75 liters the other will be 45 liters